# Retraining with Predicted Hard Labels Provably Increases Model Accuracy

### v1

@@ Appendix E Proof of Theorem 4
$\mathbb{E}\Big{[}\exp\big{(}r\widetilde{\mathrm{y}}_{1}^{\prime}\langle{\mathbf{x}}_{1}^{\prime},{\mathbf{x}}\rangle\big{)}\Big{]}\leq\exp\big{(}\mu\sigma r{\mathrm{x}}^{(1)}\big{)}{\exp\Bigg{(}\frac{(\sigma r\|{\mathbf{x}}\|)^{2}}{2}\Bigg{)}}+p_{2}\sqrt{h(r)}.$ (118)
Let us pick $r=r^{*}:=-\frac{\mu{\mathrm{x}}^{(1)}(1-2p_{2})}{2\sigma\|{\mathbf{x}}\|^{2}}$ (recall ${\mathrm{x}}^{(1)}>0$). Note that $r^{*}<0$ and we needed $r$ to be $<0$ for the Chernoff bound in eq. 105 to hold. With this value of $r$, we have:
$\exp\big{(}\mu\sigma r^{*}{\mathrm{x}}^{(1)}\big{)}{\exp\Bigg{(}\frac{(\sigma r^{*}\|{\mathbf{x}}\|)^{2}}{2}\Bigg{)}}$ $=\exp\Bigg{(}-\frac{(\mu{\mathrm{x}}^{(1)})^{2}(1-2p_{2})}{2\|{\mathbf{x}}\|^{2}}\Bigg{(}1-\frac{1-2p_{2}}{4}\Bigg{)}\Bigg{)}$
$\leq\exp\Bigg{(}-\frac{3(\mu{\mathrm{x}}^{(1)})^{2}(1-2p_{2})}{8\|{\mathbf{x}}\|^{2}}\Bigg{)}.$ (119)
Let $u:=\frac{(\mu{\mathrm{x}}^{(1)})^{2}}{\|{\mathbf{x}}\|^{2}}$. The above equation in terms of $u$ is:
$\exp\big{(}\mu\sigma r^{*}{\mathrm{x}}^{(1)}\big{)}{\exp\Bigg{(}\frac{(\sigma r^{*}\|{\mathbf{x}}\|)^{2}}{2}\Bigg{)}}\leq\exp\Big{(}-\frac{3u}{8}(1-2p_{2})\Big{)}.$ (120)
Also, recall that:
$h(r^{*})$ $=2\text{cosh}\big{(}2\mu\sigma r^{*}{\mathrm{x}}^{(1)}\big{)}\exp\big{(}2{(\sigma r^{*}\|{\mathbf{x}}\|)^{2}}\big{)}-2.$
It can be checked that:
$h(r^{*})={2\text{cosh}\big{(}u(1-2p_{2})\big{)}\exp\Big{(}\frac{u}{2}(1-2p_{2})^{2}\Big{)}-2}.$ (121)
Since $\mu\leq\frac{1}{2}$ and ${\mathrm{x}}^{(1)}\leq\|{\mathbf{x}}\|$, $u\leq\frac{1}{4}$ and thus, $u(1-2p_{2})\leq\frac{1}{4}$ and $\frac{u}{2}(1-2p_{2})^{2}\leq\frac{1}{8}$. Using Lemma 8 and Lemma 9 above, we get:
$h(r^{*})$ $\leq 2\Big{(}1+\frac{8}{15}u^{2}(1-2p_{2})^{2}\Big{)}\Big{(}1+\frac{4}{7}u(1-2p_{2})^{2}\Big{)}-2$ (122)
$={8}u(1-2p_{2})^{2}\Bigg{(}\frac{1}{7}+\frac{2}{15}u+\frac{8}{105}u^{2}(1-2p_{2})^{2}\Bigg{)}.$ (123)
Further, using the fact that $u\leq\frac{1}{4}$ and $u^{2}(1-2p_{2})^{2}\leq\frac{1}{16}$ above, we get:
$h(r^{*})\leq{2u}(1-2p_{2})^{2}.$ (124)
Plugging in eq. 120 and eq. 124 into eq. 118, we get for $r=r^{*}$:
@@
$\mathbb{E}\Big{[}\exp\big{(}r^{*}\widetilde{\mathrm{y}}_{1}^{\prime}\langle{\mathbf{x}}_{1}^{\prime},{\mathbf{x}}\rangle\big{)}\Big{]}$ $\leq\exp\Big{(}-\frac{3u}{8}(1-2p_{2})\Big{)}+{p_{2}\sqrt{2u}}(1-2p_{2})$
$=\exp\Big{(}-\frac{3u}{8}(1-2p_{2})\Big{)}\Bigg{(}1+{p_{2}\sqrt{2u}}(1-2p_{2})\exp\Big{(}\frac{3u}{8}(1-2p_{2})\Big{)}\Bigg{)}.$
Now note that $\exp\Big{(}\frac{3u}{8}(1-2p_{2})\Big{)}\leq e^{\frac{3}{32}}$ as $u(1-2p_{2})\leq\frac{1}{4}$. Using this above together with the fact that $\sqrt{2}\times e^{\frac{3}{32}}\leq 2$, we get:
$\mathbb{E}\Big{[}\exp\big{(}r^{*}\widetilde{\mathrm{y}}_{1}^{\prime}\langle{\mathbf{x}}_{1}^{\prime},{\mathbf{x}}\rangle\big{)}\Big{]}$ $\leq\exp\Big{(}-\frac{3u}{8}(1-2p_{2})\Big{)}\Big{(}1+2{p_{2}\sqrt{u}}(1-2p_{2})\Big{)}$
$\leq\exp\Big{(}-\frac{3u}{8}(1-2p_{2})+2{p_{2}\sqrt{u}}(1-2p_{2})\Big{)},$ (125)
where the last step follows by using the fact that $\exp(z)\geq 1+z$ for all $z\in\mathbb{R}$. Plugging in the value of $u=\frac{(\mu{\mathrm{x}}^{(1)})^{2}}{\|{\mathbf{x}}\|^{2}}$ above and recalling that $\big{|}\textup{cos}\big{(}\beta({\mathbf{x}})\big{)}\big{|}=\frac{|{\mathrm{x}}^{(1)}|}{\|{\mathbf{x}}\|}$, we get:
$\mathbb{E}\Big{[}\exp\big{(}r^{*}\widetilde{\mathrm{y}}_{1}^{\prime}\langle{\mathbf{x}}_{1}^{\prime},{\mathbf{x}}\rangle\big{)}\Big{]}$ $\leq\exp\Bigg{(}-(1-2p_{2})\Bigg{(}\frac{3\mu^{2}}{8}\textup{cos}^{2}\big{(}\beta({\mathbf{x}})\big{)}-{2p_{2}\mu}\big{|}\textup{cos}\big{(}\beta({\mathbf{x}})\big{)}\big{|}\Bigg{)}\Bigg{)}.$ (126)
Plugging in eq. 126 into eq. 107 and then applying that into eq. 106 yields:
$\mathbb{E}\Big{[}\exp\big{(}r^{*}\varepsilon^{\prime}\big{)}\Big{]}\leq\exp\Bigg{(}-n(1-2p_{2})\Bigg{(}\frac{3\mu^{2}}{8}\textup{cos}^{2}\big{(}\beta({\mathbf{x}})\big{)}-{2p_{2}\mu}\big{|}\textup{cos}\big{(}\beta({\mathbf{x}})\big{)}\big{|}\Bigg{)}\Bigg{)}.$ (127)
Next, plugging this into eq. 105 gives us:
$\mathbb{P}\big{(}\varepsilon^{\prime}\leq 0\big{)}\leq\exp\Bigg{(}-n(1-2p_{2})\Bigg{(}\underbrace{\frac{3\mu^{2}}{8}\textup{cos}^{2}\big{(}\beta({\mathbf{x}})\big{)}-{2p_{2}\mu}\big{|}\textup{cos}\big{(}\beta({\mathbf{x}})\big{)}\big{|}}_{:=\text{(II)}}\Bigg{)}\Bigg{)}.$ (128)
[…]
$\mathbb{P}\big{(}\varepsilon^{\prime}\leq 0\big{)}\leq\exp\Bigg{(}-\frac{n\mu^{2}}{4}(1-2p_{2})\textup{cos}^{2}\big{(}\beta({\mathbf{x}})\big{)}\Bigg{)}.$ (129)
Recalling the definition of $\varepsilon^{\prime}$ (eq. 104) and using the result above that in eq. 103, we obtain the following bound for the case of ${\mathrm{y}}=1$:
@@
$\mathbb{P}\Big{(}\langle{\mathbf{x}},\widehat{\bm{\theta}}_{1}\rangle\leq 0\Big{)}\leq\exp\Bigg{(}-\frac{n\mu^{2}}{4}(1-2p_{2})\textup{cos}^{2}\big{(}\beta({\mathbf{x}})\big{)}\Bigg{)},$ (130)
as long as $|\textup{cos}\big{(}\beta({\mathbf{x}})\big{)}|\geq\frac{{16p_{2}}}{\mu}$ with $p_{2}=\exp\Big{(}-\frac{9(d-1)}{40}\Big{)}+\exp\Big{(}-\frac{n\mu^{2}}{12d}(1-2p)^{2}\gamma^{2}\Big{)}$. This finishes the proof of the result for the case of ${\mathrm{y}}=1$. The result for the other case, i.e. ${\mathrm{y}}=-1$, can be derived similarly as above by analyzing with $r>0$ instead. In particular, for the case of ${\mathrm{y}}=-1$, we get:
@@
$\mathbb{P}\Big{(}\langle{\mathbf{x}},\widehat{\bm{\theta}}_{1}\rangle>0\Big{)}\leq\exp\Bigg{(}-\frac{n\mu^{2}}{4}(1-2p_{2})\textup{cos}^{2}\big{(}\beta({\mathbf{x}})\big{)}\Bigg{)},$ (131)
again as long as $|\textup{cos}\big{(}\beta({\mathbf{x}})\big{)}|\geq\frac{{16p_{2}}}{\mu}$. ∎
Lemma 8.
For $t\leq\frac{1}{4}$, it holds that:
$\textup{cosh}(t)\leq 1+\frac{8t^{2}}{15}.$
Proof.
Using the Taylor expansion of $\textup{cosh}(t)$, we get:
$\textup{cosh}(t)$ $=1+\frac{t^{2}}{2!}+\frac{t^{4}}{4!}+\ldots$ (132)
$\leq 1+\frac{t^{2}}{2}\Big{(}1+t^{2}+\ldots\Big{)}$ (133)
$=1+\frac{t^{2}}{2(1-t^{2})}$ (134)
Since $t\leq\frac{1}{4}$, $1-t^{2}\geq\frac{15}{16}$. Using this above gives us the desired result. ∎
Lemma 9.
For $t\leq\frac{1}{8}$, it holds that:
$e^{t}\leq 1+\frac{8t}{7}.$
Proof.
Using the Taylor expansion of $e^{t}$, we get:
$e^{t}$ $=1+t+\frac{t^{2}}{2!}+\frac{t^{3}}{3!}+\frac{t^{4}}{4!}+\ldots$ (135)
$\leq 1+t\big{(}1+t+t^{2}+\ldots\big{)}$ (136)
$=1+\frac{t}{(1-t)}.$ (137)
Since $t\leq\frac{1}{8}$, $(1-t)\geq\frac{7}{8}$. Using this above gives us the desired result. ∎
Appendix F Proof of Theorem 5
Proof.
Just like the proof of Theorem 2, we can prove the result for $\bm{\theta}^{\ast}={\mathbf{e}}_{1}$ and =$\mathcal{B}=\{{\mathbf{e}}_{1},{\mathbf{e}}_{2},\ldots,{\mathbf{e}}_{d}\}$ without loss of generality. In that case, ${\mathrm{a}}^{(j)}={\mathrm{x}}^{(j)}=\langle{\mathbf{x}},{\mathbf{e}}_{j}\rangle$ for $j\in[d]$, the ground truth label of ${\mathbf{x}}$ is $\text{sign}({\mathrm{x}}^{(1)})$ and $\big{|}\textup{cos}\big{(}\beta({\mathbf{x}})\big{)}\big{|}=\frac{|{\mathrm{x}}^{(1)}|}{\|{\mathbf{x}}\|}$.
For the above case, we have: [^14] To be explicit, in the series of equations below, we can go from eq. 140 to eq. 141 because $\mathcal{X}_{L}\cup\mathcal{X}_{U}$ has zero measure.
$\text{acc}(\widehat{\bm{\theta}}_{1})$ $=\mathbb{P}_{{\mathbf{x}}\sim\mathcal{D},\mathcal{T},\mathcal{T}_{2}}\Big{(}\textup{sign}\big{(}\langle{\mathbf{x}},\widehat{\bm{\theta}}_{1}\rangle\big{)}=\text{sign}({\mathrm{x}}^{(1)})\Big{)}$ (138)
$=\mathbb{E}_{{\mathbf{x}}\sim\mathcal{D},\mathcal{T},\mathcal{T}_{2}}\Bigg{[}\mathbbm{1}\Big{(}\textup{sign}\big{(}\langle{\mathbf{x}},\widehat{\bm{\theta}}_{1}\rangle\big{)}=\textup{sign}\big{(}{\mathrm{x}}^{(1)}\big{)}\Big{)}\Bigg{]}$ (139)
$=\mathbb{E}_{{\mathbf{x}}\sim\mathcal{D}}\Bigg{[}\mathbb{P}_{\mathcal{T},\mathcal{T}_{2}}\Big{(}\textup{sign}\big{(}\langle{\mathbf{x}},\widehat{\bm{\theta}}_{1}\rangle\big{)}=\textup{sign}\big{(}{\mathrm{x}}^{(1)}\big{)}\Big{)}\Bigg{]}$ (140)
$\geq\mathbb{E}_{{\mathbf{x}}\sim\mathcal{D}}\Bigg{[}\alpha_{1}({\mathbf{x}})\Big{|}|\textup{cos}\big{(}\beta({\mathbf{x}})\big{)}|\geq\frac{\gamma}{\sqrt{\gamma^{2}+{3}(d-1)}}\Bigg{]}\mathbb{P}\Bigg{(}|\textup{cos}\big{(}\beta({\mathbf{x}})\big{)}|\geq\frac{\gamma}{\sqrt{\gamma^{2}+{3}(d-1)}}\Bigg{)},$ (141)
where $\alpha_{1}({\mathbf{x}})=1-\exp\Big{(}-\frac{n\mu^{2}}{4}(1-2p_{2})\textup{cos}^{2}\big{(}\beta({\mathbf{x}})\big{)}\Big{)}$ is as defined in Theorem 4 and we will ensure $\frac{\gamma}{\sqrt{\gamma^{2}+{3}(d-1)}}\geq\frac{{16p_{2}}}{\mu}$, which can be simplified to:
$p_{2}\leq\Theta\Big{(}\frac{\gamma\mu}{\sqrt{d}}\Big{)}.$
Recall that $p_{2}=\exp\Big{(}-\frac{9(d-1)}{40}\Big{)}+\exp\Big{(}-\frac{n\mu^{2}\gamma^{2}}{12d}(1-2p)^{2}\Big{)}=\exp\big{(}-\Theta(d)\big{)}+\exp\Big{(}-\frac{n\mu^{2}\gamma^{2}(1-2p)^{2}}{\Theta(d)}\Big{)}$. Note that $\exp\big{(}-\Theta(d)\big{)}\leq\Theta\Big{(}\frac{\gamma\mu}{\sqrt{d}}\Big{)}$ when $\gamma\geq\Theta\Big{(}\frac{\sqrt{d}\exp(-\Theta(d))}{\mu}\Big{)}$. Imposing $\exp\Big{(}-\frac{n\mu^{2}\gamma^{2}(1-2p)^{2}}{\Theta(d)}\Big{)}\leq\Theta\Big{(}\frac{\gamma\mu}{\sqrt{d}}\Big{)}$ gives us:
$n\geq\Theta\Bigg{(}\frac{1}{(1-2p)^{2}}\Big{(}\frac{d}{\mu^{2}\gamma^{2}}\Big{)}\log\Big{(}\frac{d}{\mu^{2}\gamma^{2}}\Big{)}\Bigg{)}.$ (142)
[…]
$n\geq\Theta\Bigg{(}\frac{1}{(1-2p)^{2}}\Big{(}\frac{d}{\mu^{2}\gamma^{2}}\Big{)}\max\Bigg{(}\log\Big{(}\frac{d}{\mu^{2}\gamma^{2}}\Big{)},\log\frac{1}{p}\Bigg{)}\Bigg{)}.$ (143)
So our final bound for $p_{2}$ is:
$p_{2}\leq\min\Big{(}p,\Theta\Big{(}\frac{\gamma\mu}{\sqrt{d}}\Big{)}\Big{)}.$ (144)
With the above constraint, we get after applying Lemma 3:
$\mathbb{P}\Bigg{(}|\textup{cos}\big{(}\beta({\mathbf{x}})\big{)}|\geq\frac{\gamma}{\sqrt{\gamma^{2}+{3}(d-1)}}\Bigg{)}\geq 1-\exp\big{(}-\Theta(d)\big{)}.$ (145)
Moreover, with our constraint:
$\mathbb{E}_{{\mathbf{x}}\sim\mathcal{D}}\Bigg{[}\alpha_{1}({\mathbf{x}})\Big{|}|\textup{cos}\big{(}\beta({\mathbf{x}})\big{)}|\geq\frac{\gamma}{\sqrt{\gamma^{2}+{3}(d-1)}}\Bigg{]}\geq 1-\exp\Bigg{(}-\frac{n{\mu^{2}\gamma^{2}}(1-2p_{2})}{\Theta(d)}\Bigg{)}.$ (146)
Plugging in eq. 145 and eq. 146 into eq. 141 gives us:
$\text{acc}(\widehat{\bm{\theta}}_{1})\geq 1-\exp\big{(}-\Theta(d)\big{)}-\exp\Bigg{(}-\frac{n{\mu^{2}\gamma^{2}}(1-2p_{2})}{\Theta(d)}\Bigg{)}.$ (147)
This completes the proof. ∎
Appendix G Roadmap of Analysis with Exact Minimizer $\bar{\bm{\theta}}_{0}$
Let
${\mathbf{P}}:=\Bigg{(}\frac{1}{n}\sum_{k=1}^{n}{\mathbf{x}}_{k}{\mathbf{x}}_{k}^{\top}-\sigma^{2}{\mathbf{I}}_{d}\Bigg{)}.$ (148)
Then, we can write:
$\bar{\bm{\theta}}_{0}=\Bigg{(}\frac{1}{n}\sum_{k=1}^{n}{\mathbf{x}}_{k}{\mathbf{x}}_{k}^{\top}\Bigg{)}^{-1}\Bigg{(}\frac{1}{n}\sum_{j=1}^{n}\widehat{\mathrm{y}}_{j}{\mathbf{x}}_{j}\Bigg{)}=\frac{1}{\sigma^{2}}\Bigg{(}{\mathbf{I}}_{d}+\frac{{\mathbf{P}}}{\sigma^{2}}\Bigg{)}^{-1}\Bigg{(}\frac{1}{n}\sum_{j=1}^{n}\widehat{\mathrm{y}}_{j}{\mathbf{x}}_{j}\Bigg{)}.$ (149)
We can bound $\|{\mathbf{P}}\|$ with high probability. For instance, using Theorem 13.3 of Rinaldo and Neopane, (2018), we have with a probability of at least $1-\delta$:
$\|{\mathbf{P}}\|\leq C\sigma^{2}\max\Bigg{(}\sqrt{\frac{d+\log(2/\delta)}{n}},\frac{d+\log(2/\delta)}{n}\Bigg{)},$ (150)
for some absolute constant $C>0$. We will focus on the case of $n>\mathcal{O}\big{(}d+\log(2/\delta)\big{)}$; in that case, we have:
$\|{\mathbf{P}}\|\leq C\sigma^{2}\sqrt{\frac{d+\log(2/\delta)}{n}},$ (151)
with a probability of at least $1-\delta$. From eq. 2, recall:
$\widehat{\bm{\theta}}_{0}=\frac{1}{\sigma^{2}}\Bigg{(}\frac{1}{n}\sum_{j=1}^{n}\widehat{\mathrm{y}}_{j}{\mathbf{x}}_{j}\Bigg{)},$ (152)
Then using Lemma 10, we get:
$\|\bar{\bm{\theta}}_{0}-\widehat{\bm{\theta}}_{0}\|\leq\Bigg{(}\frac{\frac{\|{\mathbf{P}}\|}{\sigma^{2}}}{1-\frac{\|{\mathbf{P}}\|}{\sigma^{2}}}\Bigg{)}\|\widehat{\bm{\theta}}_{0}\|.$ (153)
Next, using eq. 151 above, we get:
$\|\bar{\bm{\theta}}_{0}-\widehat{\bm{\theta}}_{0}\|\leq\Bigg{(}\frac{C\sqrt{\frac{d+\log(2/\delta)}{n}}}{1-{C}\sqrt{\frac{d+\log(2/\delta)}{n}}}\Bigg{)}\|\widehat{\bm{\theta}}_{0}\|,$ (154)
with a probability of at least $1-\delta$. Note that:
@@ Appendix G Roadmap of Analysis with Exact Minimizer $\bar{\bm{\theta}}_{0}$
$\langle{\mathbf{x}}^{\prime},\bar{\bm{\theta}}_{0}\rangle$ $=\langle{\mathbf{x}}^{\prime},\widehat{\bm{\theta}}_{0}\rangle+\langle{\mathbf{x}}^{\prime},\bar{\bm{\theta}}_{0}-\widehat{\bm{\theta}}_{0}\rangle.$
Using the Cauchy-Schwarz inequality and eq. 154 above, we get:

### v2

@@ B.2 Lower bound $\widetilde{\alpha}_{0}(\mathbf{x})$
$\operatorname{\mathbb{E}}\left[\exp\left(-\frac{n(1-2q)^{2}\langle y\bm{x},{\bm{{\mu}}}\rangle^{2}}{64(\langle\bm{x},{\bm{{\mu}}}\rangle^{2}+\left\|\bm{\Sigma}^{1/2}\bm{x}\right\|_{\ell_{2}}^{2})}\right)\mathbb{1}_{\mathcal{E}_{0}}\right]$ $\leq\operatorname{\mathbb{E}}\left[\exp\left(-\frac{n(1-2q^{\prime})^{2}\langle y\bm{x},{\bm{{\mu}}}\rangle^{2}}{64(\langle\bm{x},{\bm{{\mu}}}\rangle^{2}+\left\|\bm{\Sigma}^{1/2}\bm{x}\right\|_{\ell_{2}}^{2})}\right)\mathbb{1}_{\mathcal{E}_{0}}\right]$
$\leq\exp\left(-\frac{n(1-2q^{\prime})^{2}\langle y\bm{x},{\bm{{\mu}}}\rangle^{2}}{64(\langle\bm{x},{\bm{{\mu}}}\rangle^{2}+\left\|\bm{\Sigma}^{1/2}\bm{x}\right\|_{\ell_{2}}^{2})}\right)\,.$
Using the above bound along with the bound on $\mathbb{P}(\mathcal{E}_{0}^{c})$ into (E.9), we arrive at
@@
$\mathbb{P}(\sum_{\ell}T_{\ell}<0)\leq 2\exp\left(-\frac{n(1-2q^{\prime})^{2}\langle y\bm{x},{\bm{{\mu}}}\rangle^{2}}{64(\langle\bm{x},{\bm{{\mu}}}\rangle^{2}+\left\|\bm{\Sigma}^{1/2}\bm{x}\right\|_{\ell_{2}}^{2})}\right)+4\exp\left(-\frac{n}{32}(1-2p)^{2}\right)\,.$ (E.10)
Invoking (E.5) and the definition of $T_{\ell}$ given by (E.6) we obtain
@@
$\mathbb{P}(y\langle\bm{x},{\widehat{\bm{\theta}}}_{1}\rangle<0;\mathcal{E})\leq 2\exp\left(-\frac{n(1-2q^{\prime})^{2}\langle y\bm{x},{\bm{{\mu}}}\rangle^{2}}{64(\langle\bm{x},{\bm{{\mu}}}\rangle^{2}+\left\|\bm{\Sigma}^{1/2}\bm{x}\right\|_{\ell_{2}}^{2})}\right)+4\exp\left(-\frac{n}{32}(1-2p)^{2}\right)\,.$ (E.11)
For the final step we combine (E.11), (E.4) and (E.3) to get
@@
$\mathbb{P}(y\langle\bm{x},{\widehat{\bm{\theta}}}_{1}\rangle<0)\leq$ $2\exp\left(-\frac{n(1-2q^{\prime})^{2}\langle y\bm{x},{\bm{{\mu}}}\rangle^{2}}{64(\langle\bm{x},{\bm{{\mu}}}\rangle^{2}+\left\|\bm{\Sigma}^{1/2}\bm{x}\right\|_{\ell_{2}}^{2})}\right)$
$+4\exp\left(-\frac{n}{32}(1-2p)^{2}\right)+\frac{n}{2}\left(\exp\left(-\frac{\gamma^{4}}{8\lambda_{\max}^{2}}(1-2p^{\prime})\frac{n}{d}\right)+e^{-d/16}\right)e^{d/n}\,.$ (E.12)
E.1 Proof of Lemma E.1
Define $\tilde{\xi}_{i}:=\xi_{i}(1+u_{i})$. By invoking (E.1) we have
@@
$\widetilde{y}_{\ell}^{\prime}\langle\bm{x}_{\ell},{\widehat{\bm{\theta}}}_{0}\rangle=\frac{\widetilde{y}_{\ell}^{\prime}}{n}\langle\bm{x}_{\ell},(\sum_{i}\tilde{\xi}_{i}){\bm{{\mu}}}+\bm{\Sigma}^{1/2}\xi_{\ell}y_{\ell}{\bm{z}}_{\ell}+\bm{\Sigma}^{1/2}\sum_{i\neq\ell}\xi_{i}y_{i}{\bm{z}}_{i}\rangle.$
We have $\sum_{i\neq\ell}\xi_{i}y_{i}{\bm{z}}_{i}\sim{\mathcal{N}}(\bm{0},(n-1)\bm{I}_{d})$ and so $\langle\widetilde{y}_{\ell}^{\prime}\bm{x}_{\ell},\bm{\Sigma}^{1/2}\sum_{i\neq\ell}\xi_{i}y_{i}{\bm{z}}_{i}\rangle$ is a zero mean Gaussian with variance $(n-1)\left\|\bm{\Sigma}^{1/2}\bm{x}_{\ell}\right\|_{\ell_{2}}^{2}$. Let $\mathcal{F}_{\ell}$ be the $\sigma$-algebra generated by $({\bm{z}}_{\ell},y_{\ell},\{\xi_{i},u_{i}\}_{i\in[n]})$. Note that ${\bm{z}}_{i}$, for $i\neq\ell$, is independent of $\mathcal{F}_{\ell}$ and $\widetilde{y}^{\prime}_{\ell}$ is $\mathcal{F}_{\ell}$-measurable. Hence,